OpenStudy (anonymous):

The equation, 6x + 3y = 15, has a slope of ______.

OpenStudy (anonymous):

not telling you haha

OpenStudy (anonymous):

6x+3y=15 6x + 3(9)=15 3*9 6x+27=15 6x=15-27 since 27 is possitive + infront..change to negative 6x=-12 x=-2 answer.. very simple problem...(-2,9)